OpenStudy (anonymous):

9. A gaseous organic compound, X, was burnt in an excess of oxygen. A 0.112 dm sample of X, measured at s.t.p., produced 0.88 g of carbon dioxide. How many carbon atoms are there in one molecule of X?

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@JFraser @Luigi0210 @Nurali

OpenStudy (anonymous):

its .0112 dm 3

OpenStudy (nurali):

please delete all name.

OpenStudy (anonymous):

why ? I just tag them so they can help me ..Not the Answer ..

OpenStudy (jfraser):

if you know how many grams of \(CO_2\) are produced, then you know how many moles of \(CO_2\) are produced. if you know how many moles of \(CO_2\) are produced, then you know how many molecules of \(CO_2\) are produced. if you know how many molecules of \(CO_2\) are produced, each molecule of \(CO_2\) must contain ONE atom of C. Find the proper ratios

OpenStudy (nurali):

First, you need to determine how many moles of the gaseous compound you began with. Since you are at STP, you know that 1 mole of any gas at STP occupies a volume of 22.4 dm^3. So, 0.112 dm^3 / 22.4 dm^3/mol = 5.00X10^-3 moles of gas X Now, from the mass of CO2 formed, you can calculate the moles of CO2 (and thus the moles of C in the compound: 0.88 g CO2 / 44.0 g/mol = 0.02 moles CO2 = 0.02 moles C Dividing moles of C by the moles of the original compound will give you the number of carbon atoms per molecule: 0.02 / 0.005 = 4. So, the compound contains 4 carbon atoms per molecule.

OpenStudy (anonymous):

I need help .not the whole question solved by you ..you could help me by telling me step by step procedure @Nurali

OpenStudy (anonymous):

Thank you so much @Nurali and @JFraser

OpenStudy (nurali):

Anytime.